Given any n∈N, is possible to define a linear transformation whose kernel has dimension n.
True or false with full explanation
1
Expert's answer
2021-08-16T15:29:12-0400

Consider the map "T:\R^n\to\R^n, T(x_1,x_2,...,x_n)=(0,0,...,0)" for any "(x_1,x_2,...,x_n)\in\R^n." Then

"T(a(x_1,x_2,...,x_n)+b(y_1,y_2,...,y_n))=0=\\=a\cdot 0+b\cdot 0=a\cdot T(x_1,x_2,...,x_n)+b\cdot T(y_1,y_2,...,y_n)," and hence "T" is a linear transformation. Its kernel is

"ker (T)=\{(x_1,x_2,..,x_n)\in\R^n:T(x_1,x_2,..,x_n)=(0,0,...,0)\}=R^n."

Since "dim(\R^n)=n," we conclude for any "n∈\N", is possible to define a linear transformation whose kernel has dimension "n" .


Answer: true